OpenStudy (insa):

Find the equation of the straight line with gradient -2/3 and passing through(-3,5). If this line also passes through the point (a,3), find a.

OpenStudy (midhun.madhu1987):

The equation of a line having gradient and passing through a point is given by y-y1 = m(x - x1) where m = Gradient/Slope Just plug in the values...

OpenStudy (midhun.madhu1987):